OpenStudy (anonymous):

determine if 1,4 is a solution of the system 6x + 9y = 42, 5x + 0y = 5

Parth (parthkohli):

Substitute x with 1 and y with 4. Solve for both and see if the equations are satisfied.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 6(1) + 9(4) = 42 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6 + 36 = 42 }\) \(\Large \color{MidnightBlue}{\Rightarrow 42 = 42 }\) First equation \(\checkmark \) \(\Large \color{MidnightBlue}{\Rightarrow 5(1) + 0(4) = 5 }\) \(\Large \color{MidnightBlue}{\Rightarrow 5 + 0 = 5}\) \(\Large \color{MidnightBlue}{\Rightarrow 5 = 5 }\) Second equation ---> \(\checkmark\) So it is a solution.

OpenStudy (anonymous):

Thanks! This is the solution.